a motorcyclist travels 90 miles at rate of 20 miles per hour, if he returns same distance at rate of 40 miles per hour, what is the average speed for entire trip, in miles pr hour?
Answer by mananth(16949)

while going
90-------------------20--------------------time = 90/20 =4.5 hours

90------------------40---------------------time = 90/40 = 2.25 hours
total time = 6.75 hours
total distance = 180 miles
Average speed = 180/6.75
= 26.7 mph